(a) The equivalent resistance of a group of resistors connected in parallel is

(i) greater than any of the resistors in the group,
(ii) less than any of the resistors in the group, or
(iii) neither of the above.

(b) The equivalent resistance of a group of resistors connected in series is

(i) greater than any of the resistors in the group,
(ii) less than any of the resistors in the group, or
(iii) neither of the above.

Justify your answers by considering a circuit consisting of a 3.0 ohm resistor and a 5.0 ohm resistor connected across a 12 v battery.
Answer #1
(a)
(ii) less than any of the resistors in the group
Equivalent resistance R = R1R2/(R1 + R2) = 15/8 = 1.875 Ω
(b)
(i) greater than any of the resistors in the group
Equivalent resistance R = R1 + R2 = 3 + 5 = 8 Ω